OpenStudy (anonymous):

write the answer in slope intercept and standard form. parallel to y=-5 and through (2,7)

OpenStudy (anonymous):

slop intercept form is y=mx+b where m is the slop and b is the y-intercept

OpenStudy (anonymous):

right. but how do i solve for the slope?

OpenStudy (anonymous):

slope formula is \[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

so: 2 is x1 and 7 is y1. but what is x2 and y2?

OpenStudy (anonymous):

y-7=slope(x-2) ? or would -5 have to be put somewhere?

OpenStudy (anonymous):

The "slope" is the gradient, that is, m in the equation y=mx +b. However, if you look at y=-5, you should see that there is no x value in that equation, hence m must be equal to 0, which means it's just a flat line. Considering the line you're looking for is parallel to this flat line, it must also be a flat line, such as y = 1, y = 2, y = 3, etc. The x-value of the point is irrelevant - consider the y-value of the point you're given, use that as the "b" in your equation, and from there you should be able to work our the new line's formula.

OpenStudy (anonymous):

If standard form means general form (we learnt different terminology) then after you've found the slope-intercept form, rearrange the equation so you have everything on one side being equal to 0.

OpenStudy (anonymous):

i think i understand. can you wait while i try to solve your equation?

OpenStudy (anonymous):

y=-5x-17

OpenStudy (anonymous):

or, to put it into y=mx+b........ y=-5x + -17

OpenStudy (anonymous):

so standard form is 5x+1y=-17, right?

OpenStudy (anonymous):

Think about it this way - there is no x in the equation, because m (the number before x) is always equal to 0, and anything multiplied by 0 is 0. What would you get then?

OpenStudy (shamim):

ok u know \[b=0\]

OpenStudy (shamim):

\[x=-2,y=7\]

OpenStudy (anonymous):

Actually I disagree, I think b = -17

OpenStudy (anonymous):

-7, I mean*

OpenStudy (shamim):

ur equation of line is\[y=mx+b\]

OpenStudy (anonymous):

Hold on, just 7... ahh I keep reading it wrong.

OpenStudy (anonymous):

I think the result would be y = 7, as that line is parallel to y = -5

OpenStudy (anonymous):

this is how i got it: y-7=-5(x- -2) (adding 7 to both sides) 7+y-7=-5x-10+7 y=-5+ -17 slope intercept: y=-5x+ -17 standard: 5x+1y=-17

OpenStudy (anonymous):

But you made an error in your working, Amenah. See, you placed -5 as the gradient in that equation, where in fact the gradient is 0. If two lines are parallel, they have the same gradient, and considering the gradient of y = -5 is 0, the gradient of the equation you're looking for must also be 0.

OpenStudy (anonymous):

so: b=-17?

OpenStudy (shamim):

y=7

OpenStudy (anonymous):

7=-5x+0 ? b=0?

OpenStudy (anonymous):

sorry: 0=-5x+7 ?

OpenStudy (anonymous):

could i send a message to you tomorrow to continue helping me? i'm being forced to get in the bed now by my mom :) Thanks for the help!

OpenStudy (anonymous):

Sure, no problem